OpenStudy (anonymous):

6x^2 + x - 12 How to factor by splitting middle term?

OpenStudy (jdoe0001):

try \( \bf 6x^2 + \color{red}{x} - 12\implies 6x^2 \color{red}{+ 9x-8x} - 12\implies (6x^2 + 9x)-(8x + 12)\)

OpenStudy (anonymous):

Man do u know how to complete the square?

OpenStudy (anonymous):

I will teach you how to solve any quadratic equation without having to factor anything: \[ax^2+bx+c=0\] It is called completing the square: Lets do it on your quadratic: \[6x^2+x-12=0\] Divide everything by 6 to get it in a form with no a-term: \[x^2+\frac{ 1 }{ 6 }x-2=0\] Add the opposite of the c-term of the quadratic to both sides in this case (2) not (-2) but 2: \[x^2+\frac{ 1 }{ 6 }x=2\] Add half of the b-term squared to both sides so half of one-sixth or one twelfth squared or" (1/144)": \[x^2+\frac{ 1 }{ 6 }x+\frac{ 1 }{ 144 }=2+\frac{ 1 }{ 144 }\] Factor this by doing be (x+half of the b-term)^2=whatever so for this quadratic:\[(x+\frac{ 1 }{ 12 })^2=2+\frac{ 1 }{ 144 }\] The 1/12th is half of the b-term i was talking about Take the square root of both sides: \[x+\frac{ 1 }{ 12 }=\pm\sqrt{2+\frac{ 1 }{ 144 }}\] Subtract 1/12 from both sides: \[x=\pm \sqrt{2+\frac{ 1 }{ 144 }}-\frac{ 1 }{ 12 }\]